OpenStudy (anonymous):

Paul drives his delivery van 675 miles in 3 days. At this rate, how far will he drive in 15 days? A. 45 miles B. 225 miles C. 2,025 miles D. 3,375 miles

OpenStudy (anonymous):

@robtobey @texaschic101 @deshawn1 @Jorg3m @Mateaus @vera_ewing

OpenStudy (anonymous):

anyone help

OpenStudy (anonymous):

Solve the following for x. \[\frac{675}{3}=\frac{x}{15} \]

OpenStudy (texaschic101):

divide 675 by 3 to find out how many miles was driven per day, then multiply that by 15. or set it up as a proportion : 675/3 = x/15....(675 miles to 3 days = x miles to 15 days) cross multiply and solve for x either way, you should arrive at the same answer

OpenStudy (anonymous):

3375 ?

OpenStudy (anonymous):

Yes.

OpenStudy (texaschic101):

you got it

OpenStudy (anonymous):

thanks